OpenStudy (anonymous):

3.What is the solution to the equation? (x8)-14=-15x=

OpenStudy (goose.):

Why is there 2 equal signs?

OpenStudy (anonymous):

hold on ill screenshot the problem

OpenStudy (goose.):

Okay.

OpenStudy (goose.):

The attatchment doesn't want to open give me a minute xD

OpenStudy (anonymous):

lol i can resend it through jpg instead of png

OpenStudy (goose.):

I got it,

OpenStudy (anonymous):

ok

OpenStudy (goose.):

You wont be distributing for this one,

OpenStudy (goose.):

Okay for starters, add 14 to both sides, to isolate your variable.

OpenStudy (goose.):

Then after you've done that, subtract 8 from both sides,

OpenStudy (goose.):

So it'll be x = -15 + 14 - 8 or you could do 14 - 23 if you combined both the -15 and -8

OpenStudy (adilalvi):

(x8)-14=-15x x8-14+15x=? 23x-14=? 23x=14 x=14/23 x=0.6086957 answer

OpenStudy (goose.):

No, they accidentally typed the problem wrong, Lol. check the screen shot.

OpenStudy (adilalvi):

am i right?

OpenStudy (anonymous):

its -9

OpenStudy (anonymous):

thanks

OpenStudy (goose.):

Yes. the answer is -9.